Problem Solving Using Systems of Equations
Julie and Eric row their boat(at a constant speed) 55 miles downstream for 5 hours, helped by the current. Rowing at the same rate, the trip back against the current takes 11 hours. Find the rate of the current

Answer by khwang(438)

Assume the rowing speed is Vr and the current speed is Vw.

We have the downstream speed: Vr + Vw = 55/5 = 11 ...(1)
and the upstream speed Vr - Vw = 55/11 = 5 ...(2)
Hence, Vr = (11+5)/2 = 8 miles/hr and Vw = (11-5)/2 = 3 miles/hr

Kenny
PS: Usually, I don't answer such kind of simple question and similar
questions appeared thousand times here.
To avoid that you got terrible long solutions so I showed something
that I learned when I was in elementary school and hope you don't
bother with this kinds of question any more.